(m+5)/15-(m-5)/10=4-(2m+10)/16
1
Expert's answer
2010-11-29T18:51:40-0500
(2(m+5) - 3(m-5))/30 = 4 - (m+5)/8;
(- m + 25)/30 = (32-m + 5) / 8;
(25 - m)/ 15 = (37 - m) / 4;
100 - 4m = 37*15 - 15 m ;
11m = 555-100 = 455;
m = 455/11 = 41 5/11.